OpenStudy (anonymous):

Rakesh needs 1.71 g of sugar (\(\bf \large C_{12}H_{22}O_{11}\)) to sweeten his tea. What would be the number of carbon atoms present in his tea? \[ \large \bf a) 3.6 \times 10^{22} \] \[ \large \bf b)7.2 \times 10^{21} \] \[\large \bf c)0.05 \times 10^{23} \] \[\large \bf d)6.6 \times 10^{22}\]

OpenStudy (anonymous):

1mole of sugar=342gm 1.71gm=1.71/342moles of sugar=0.005moles In 1 mole of sugar we have 12*6*10^23 carbon atoms 0.005moles=0.005*12*6*10^23=0.36*10^23 or 3.6*10^22

OpenStudy (anonymous):

So what you need to do is convert the grams of super into moles, and then use Avogadro's number to determine the number of atoms in the molecule. \[1.71g \left( \frac{ 1 mol }{ 342.308g } \right)\left( \frac{ 6.02*10^{23} }{ 1mol } \right)\left( \frac{ 12 atoms }{ 1 molecule } \right)\]